How many gram formula units are present in 1600 grams of ‘NaCl’ sample ?
 

Grade:8

2 Answers

Bhavya
7 years ago
Dear student
Gram formula unit or gram formula mass represents moles.
Moles of NaCl = Given mass/Formula mass
= 1600/58.5
= 27.35
Siddharth Singh
6 years ago
Moles of nacl is =1600/58.5=27.34Here 1600 is given mass and 58.5 is the molecular mass of nacl beacuse na has 23 gm and cl has 35.5 gm